# Beauty and the Bias: Exploring the Impact of Attractiveness on Multimodal Large Language Models

### v2

@@ 4 Results
The influence of facial attractiveness was evident across stereotyped jobs, traits, and conditions, suggesting that MLLMs systematically rely on attractiveness as a decision-making cue. As expected, attractiveness mattered in 91.1% of the attractiveness stereotyped jobs. However, even in contexts where attractiveness provides no meaningful information, such as in the stereotyped trait scenarios denoting sentiment, it impacted outcomes in 92.6% of the cases. This indicates a pervasive and potentially unwarranted reliance on attractiveness by MLLMs across diverse decision contexts, which has been understudied in the literature to date.
@@
To investigate the attractiveness halo effect, we focus on the 33 sentiment-related scenarios from the stereotyped traits category, where the “Stereotyped Choice” represents a positive trait and the other choice is the corresponding negative trait (e.g. trustworthy vs untrustworthy). The entire list of choice pairs used can be found in Table 7 in Appendix C.
Statistically significant differences were observed (Kruskal-Wallis, $p<0.01$) in 92.6% of these scenarios on average across models. In all scenarios and for all models except for 3 out of 31 scenarios for DeepSeek and 1 out of 30 for Qwen2, the beautified images were associated with the positive traits when there was a significant difference. Hence, we find strong evidence on the existence of the attractiveness halo effect in MLLMs. These findings suggest that decisions made by MLLMs are not only influenced by facial attractiveness but also that, similarly to humans, they have a strong tendency to associate attractive individuals with positive traits.
@@
The nature of our scenarios and input data allows us to evaluate the existence and strength of gender, race, and age biases in MLLMs. We first assess these biases independently of attractiveness, and then extend our analysis by considering their intersection.
@@
1. Gender-stereotyped jobs. On average across models, 93.2% of the gender-stereotyped job scenarios exhibited statistically significant effects (Kruskal–Wallis, $p<0.01$), with different responses depending on the gender of the face. In the case of Phi3.5, it provided different responses depending on gender in all 19 gender-stereotyped job scenarios. Crucially, in every instance where a significant effect was identified, MLLMs were more likely to associate male faces with male-stereotyped jobs, replicating social gender stereotypes and hence exhibiting a gender bias. Appendix F includes a list of the gender-stereotyped job scenarios and the magnitude of the observed effects for all evaluated models.
2. Race-stereotyped jobs.
A statistically significant effect (Kruskal–Wallis, $p<0.01$) of the race of the individual depicted in the image was observed in 35.7% of the scenarios across models when testing for a racial bias between subgroups corresponding to the race stereotypes used to define the scenario (see Table 6 in Appendix C for the list of race stereotyped jobs and the corresponding races). In 93.3% of the cases where a significant effect was detected, the MLLMs were more likely to associate images in ways that conformed to prevailing social stereotypes tied to the respective job pairs as indicated in the Table. The list of race-stereotyped occupations and corresponding effect sizes across all models are provided in Appendix G.
3. Stereotyped conditions.
The most pronounced effects in this category occurred with the scenarios concerning geography-based stereotypes, particularly regarding race biases. Images of White individuals were significantly less likely to be classified as “immigrant” or “foreign” compared to other racial groups, whereas images of Indian individuals were the most likely to be classified as such. This finding aligns with prior research suggesting that large language models tend to reflect the biases of WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societies [Atari et al.(2023)]. Similarly, Black individuals were the least likely to be classified as “educated”, although this effect, while statistically significant, was relatively small. No significant effects of age or gender were observed in the education-related stereotypes. While some statistically significant effects were observed across these categories, the overall effect sizes were generally smaller than those reported for other types of biases in this analysis.
@@
We further investigate whether the impact of the attractiveness bias varies across different gender (${W}_{\textit{gender}}^{\textit{attr}}$), age (${W}_{\textit{age}}^{\textit{attr}}$), and racial groups (${W}_{\textit{race}}^{\textit{attr}}$). Thus, we compute the attractiveness bias independently for each subgroup and conduct Bonferroni-corrected Wilcoxon Paired Rank tests across the 91 scenarios for every pair of subgroups to assess whether the observed differences are statistically significant. Table 5 presents the differential strength of attractiveness bias between images of males and females. Across all evaluated models, we consistently find a stronger attractiveness bias associated with female images, in alignment with previous findings from studies with human participants [Gulati et al.(2024)].
@@ 5 Discussion
Our study provides compelling empirical evidence for the existence of an attractiveness bias influencing judgments made by MLLMs. While attractiveness is hard to study due to its highly subjective nature, our methodology relies on beauty filters that increase the attractiveness of individuals without impacting their identity thus minimizing confounds. Complementary studies involving human participants rating the images used in our study [Gulati et al.(2024)] confirmed that individuals were perceived as significantly more attractive when the beauty filter was applied, thus validating the attractiveness manipulation used here. We observed statistically significant differences in 86.2% of the scenarios where the MLLMs evaluated images of the same individuals before and after applying a beauty filter indicating that this bias is indeed prevalent in decisions made by MLLMs.
@@
Our research provides compelling evidence demonstrating that attractiveness significantly influences the MLLMs decision-making processes, potentially comparable in magnitude to traditional demographic variables. Crucially, the associations between attractiveness and positive evaluative terms occurs in a non-transparent and implicit manner. This opacity could induce naive end-users to the mistaken belief that MLLM-generated decisions are objective and unbiased. Given the accelerating adoption of these models in high-stakes scenarios, such as recruitment and professional evaluations [^6] https://www.theguardian.com/technology/2019/oct/25/unilever-saves-on-recruiters-by-using-ai-to-assess-job-interviews, the subtle yet consequential biases related to attractiveness can inadvertently lead to discriminatory outcomes or unjustified preferential treatment.
Consequently, our study underscores the urgent need for research in the design, development, and validation of MLLMs. It highlights the imperative to extend bias-mitigation strategies beyond the traditional emphasis on demographic factors, proactively addressing cognitive biases such as the attractiveness halo effect.
@@
The identified attractiveness bias does not operate in isolation but interacts with other salient societal biases, including those related to gender, age, and race. Our analysis reveals that this bias is significantly more pronounced when evaluating female subjects. This finding aligns with prior human-subject studies, which have demonstrated that the attractiveness halo effect exerts a stronger influence in the evaluation of female faces [Kunst et al.(2023), Gulati et al.(2024)]. In addition to the heightened attractiveness bias for female subjects, we also observed an amplification of gender biases in the subset of images that had been altered using beauty filters. This observation mirrors patterns identified in human assessments of beautified faces [Gulati et al.(2024)].
We also observed a relative attenuation of racial and age related biases in the beautified image sets for certain MLLMs. We hypothesize that this differential behavior with race biases in the beautified images could be due to the homogenization of the facial features across different racial groups that has been reported to occur after the application of beauty filters [Riccio and Oliver(2022), Riccio et al.(2022)]. Beauty filters were also found to make people look younger [Gulati et al.(2024)], which could explain the relative reduction of the observed age related biases for middle-aged individuals.
As previously noted, the influence of attractiveness bias varies across different demographic groups, complicating efforts toward effective bias mitigation. Prior research has shown that mitigation strategies targeting a single demographic attribute – such as gender – can unintentionally intensify biases associated with other attributes, such as race [Ramachandranpillai et al.(2024)]. These unintended cross-demographic interactions may similarly apply to biases arising from perceptions of attractiveness. Moreover, existing approaches designed to counteract gender or race based biases may inadvertently reinforce attractiveness related biases if these dimensions are not explicitly considered during model design and evaluation. Such interdependencies underscore the complexity of bias mitigation in MLLMs and highlight the need for further research into holistic, intersectional mitigation frameworks that address both traditional demographic variables and less-studied factors like attractiveness.
@@
Our work is not exempt from limitations. First, we do not evaluate API-based models such as GPT-4, as our experimental paradigm involves providing facial images to the models as inputs. Due to the lack of transparency related to data handling by API providers, and restrictions imposed by the dataset license (which specifically prohibits the use of images in such contexts) we excluded these models from our evaluation. Nevertheless, our analysis spans seven different open-source models, within which we observe statistically significant and consistent patterns, underscoring the robustness and prevalence of attractiveness-related biases. Second, while we carefully designed and curated our inputs to minimize confounding variables, our empirical evaluation may not fully capture the diversity of real-world contexts where attractiveness plays a role across cultures and social settings. Third, we have used a discrete number of categories for gender, age and race, as given by the ground truth labels of the dataset, which inevitably fails to capture and represent the diversity in society. Lastly, while we examine intersections with gender, age and race, we have not considered all relevant social factors, including socioeconomic status or disability, which could further impact the identified biases.
@@ 6 Conclusion
In this paper, we have studied the role that attractiveness plays in the decision-making processes of MLLMs by evaluating seven different models in 91 scenarios with over 900 face images and more than 7,000,000 prompts. Our findings provide strong evidence of the existence of an attractiveness bias in decisions made by MLLMs which can manifest in ways that mimic the attractiveness halo effect, a cognitive bias observed in humans. We also find evidence of the complex interplay between attractiveness and demographic factors –namely gender, age, and race– in driving the decisions of the MLLM. This interaction between attractiveness and other factors increases the complexity of interventions aimed at mitigating biases in MLLMs. We hope that not only our results, but also our dataset and controlled experimental design, will serve as tools for the research community to further measure and understand these biases in MLLMs.
Acknowledgements
@@ Appendix C Scenarios
Model Makeup artist GPT-4o
Acting professional Screen writer GPT-4o
Personal trainer Engineer GPT-4o
Real estate agent Computer Scientist GPT-4o
Salesperson Product inspector GPT-4o
News Anchor Copywriter GPT-4o
Influencer Factory worker GPT-4o
Public relations specialist Translator GPT-4o
@@ Appendix E Attractiveness Halo Effect in Stereotyped Traits
Tables 10 - 16 below detail the strength of the attractiveness halo effect observed in all tested models. Each table reports the test statistic from the Kruskal–Wallis test (${H}_{i}^{\textit{attr}}$) comparing the values of $\phi_{i}$ between the beautified and original image groups for each scenario $s_{i}$. Standard star notation is employed to indicate the level of statistical significance: ***∗ ∗ ∗ denotes $p<0.001$ and **∗ ∗ denotes $p<0.01$. Additionally, the tables provide the mean values of $p_{i}$ for both beautified and original images. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the first choice. This pattern reflects the model’s underlying preference or bias. Notably, across nearly all scenarios, the model demonstrates a consistent inclination to associate beautified images with the first choice, which corresponds to a positive sentiment (“Stereotyped Choice”). This trend underscores the presence of a robust attractiveness halo effect influencing the model’s decision-making.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 10: Attractiveness halo effect in sentiment oriented stereotyped traits for Gemma. Out of 33 scenarios, 31 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 11: Attractiveness halo effect in sentiment oriented stereotyped traits for Phi3.5. Out of 33 scenarios, 28 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 12: Attractiveness halo effect in sentiment oriented stereotyped traits for DeepSeek. Out of 33 scenarios, 28 scenarios showed a significant attractiveness halo effect, while 3 showed an attractiveness bias but in the opposite direction.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 13: Attractiveness halo effect in sentiment oriented stereotyped traits for Molmo. Out of 33 scenarios, 30 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 14: Attractiveness halo effect in sentiment oriented stereotyped traits for Qwen2. Out of 33 scenarios, 28 scenarios showed a significant attractiveness halo effect, while 1 showed an attractiveness bias but in the opposite direction.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 15: Attractiveness halo effect in sentiment oriented stereotyped traits for Pixtral. Out of 33 scenarios, 33 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 Mean $\text{Mean }\phi_{i}(x^{b})$ Mean $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 16: Attractiveness halo effect in sentiment oriented stereotyped traits for LLaVA 1.5. Out of 33 scenarios, 32 scenarios showed a significant attractiveness halo effect.

### v3

@@ 4 Results
The influence of facial attractiveness was evident across stereotyped jobs, traits, and conditions, suggesting that MLLMs systematically rely on attractiveness as a decision-making cue. As expected, attractiveness mattered in 91.1% of the attractiveness stereotyped jobs. However, even in contexts where attractiveness provides no meaningful information, it impacted a significant portion of decisions. This indicates a pervasive and potentially unwarranted reliance on attractiveness by MLLMs across diverse decision contexts, which has been understudied in the literature to date.
@@
To investigate the attractiveness halo effect, we focus on the 33 sentiment-related scenarios from the stereotyped traits category, where the “Stereotyped Choice” reflects a positive trait and the alternative represents its negative counterpart (e.g. trustworthy vs untrustworthy). The list of choice pairs used can be found in Table 7 in Appendix C.
Statistically significant differences were observed (Kruskal-Wallis, $p<0.01$) in 92.6% of these scenarios on average across models. In all scenarios and for all models (except for 3 out of 31 scenarios for DeepSeek and 1 out of 30 for Qwen2) the beautified images were associated with the positive traits and the differences with the original images were significant. This provides strong evidence of the attractiveness halo effect in MLLMs, suggesting that, like humans, these models tend to associate attractive faces with positive traits.
@@
The design of the scenarios and inputs enables the evaluation of gender, race, and age biases in MLLMs. We first assess each bias independently of attractiveness, followed by an examination of their intersection.
@@
1. Gender-stereotyped jobs. On average across models, 93.2% of the gender-stereotyped job scenarios exhibited statistically significant effects (Kruskal–Wallis, $p<0.01$), with responses varying by the gender of the face. Phi3.5 showed gender-based differences in all 19 such scenarios. Crucially, in every instance where a significant effect was identified, MLLMs were more likely to associate male faces with male-stereotyped jobs, replicating social gender stereotypes and hence exhibiting a gender bias. Appendix F details the scenarios and effect sizes across models.
2. Race-stereotyped jobs. A statistically significant effect (Kruskal–Wallis, $p<0.01$) of race was found in 35.7% of scenarios across models when comparing subgroups defined by race stereotypes (see Table 6 in Appendix C.
In 93.3% of the cases where a significant effect was detected, the MLLMs were more likely to associate images in ways that conformed to prevailing social stereotypes tied to the respective job pairs as indicated in the Table. The list of race-stereotyped occupations and corresponding effect sizes across all models is in App. G.
3. Stereotyped conditions. The most pronounced effects in this category were found in the scenarios concerning geography-based stereotypes, particularly regarding race biases. Images of White individuals were significantly less likely to be classified as “immigrant” or “foreign” compared to other racial groups, whereas images of Indian individuals were the most likely to be classified as such. This finding aligns with prior research suggesting that LLMs tend to reflect the biases of WEIRD (Western, Educated, Industrialized, Rich, and Democratic) societies [Atari et al.(2023)]. Similarly, Black individuals were the least likely to be classified as “educated”, although this effect, while statistically significant, was relatively small. No significant effects of age or gender were observed in the education-related stereotypes. While some statistically significant effects were observed across these categories, the overall effect sizes were generally smaller than those reported for other types of biases.
@@
We further investigate whether the impact of the attractiveness bias varies across different gender (${W}_{\textit{gender}}^{\textit{attr}}$), age (${W}_{\textit{age}}^{\textit{attr}}$), and racial groups (${W}_{\textit{race}}^{\textit{attr}}$). Thus, we compute the attractiveness bias independently for each subgroup and conduct Bonferroni-corrected Wilcoxon Paired Rank tests across the 91 scenarios for every pair of subgroups to assess whether the observed differences are statistically significant. Table 5 presents the differential strength of attractiveness bias between images of males and females. Across all evaluated models, we consistently find a stronger attractiveness bias associated with female images, in alignment with previous findings with human participants [Gulati et al.(2024)].
@@ 5 Discussion
Our study provides compelling empirical evidence for the existence of an attractiveness bias influencing judgments made by MLLMs. While attractiveness is hard to study due to its highly subjective nature, our methodology relies on beauty filters that increase the attractiveness of individuals without impacting their identity thus minimizing confounds. Previous studies involving human participants that rated the images used in our study [Gulati et al.(2024)] confirmed that individuals were perceived as significantly more attractive when the beauty filter was applied, thus validating the attractiveness manipulation used in this work. Statistically significant differences occurred in 86.2% of scenarios where MLLMs evaluated the same individuals before and after applying a beauty filter, indicating this bias is prevalent in MLLM decisions.
@@
Our research provides compelling evidence demonstrating that attractiveness significantly influences the decision-making processes in MLLMs, comparable in magnitude to traditional demographic variables. Crucially, the associations between attractiveness and positive traits occurs in a non-transparent and implicit manner. This opacity could induce naive end-users to the mistaken belief that MLLM-generated decisions are objective and unbiased. Given the accelerating adoption of these models in high-stakes scenarios, such as recruitment and professional evaluations [^6] https://www.theguardian.com/technology/2019/oct/25/unilever-saves-on-recruiters-by-using-ai-to-assess-job-interviews, the subtle yet consequential biases related to attractiveness can inadvertently lead to discriminatory outcomes or unjustified preferential treatment.
Thus, our study underscores the urgent need for research in the design, development, and validation of MLLMs. It underscores the need to expand bias-mitigation strategies beyond demographic factors to proactively include cognitive biases like the attractiveness halo effect.
@@
The identified attractiveness bias does not operate in isolation but interacts with other societal biases, including those related to gender, age, and race. Our analysis reveals that this bias is significantly more pronounced when evaluating females. This finding aligns with prior human-subject studies, which have demonstrated that the attractiveness halo effect exerts a stronger influence in the evaluation of female faces [Kunst et al.(2023), Gulati et al.(2024)]. In addition to the heightened attractiveness bias for female subjects, we also observed an amplification of gender biases in the subset of images that had been altered using beauty filters. This observation mirrors patterns identified in human assessments of beautified faces [Gulati et al.(2024)].
We also observed a relative attenuation of racial and age related biases in the beautified image sets for certain MLLMs. We hypothesize that this differential behavior with race biases could be due to the homogenization of the facial features across different racial groups after applying beauty filters as reported in [Riccio and Oliver(2022), Riccio et al.(2022)]. Beauty filters have also been found to make people look younger [Gulati et al.(2024)], which could explain the relative reduction of the observed age related biases for middle-aged individuals.
The influence of the attractiveness bias varies across different demographic groups, complicating efforts toward effective bias mitigation. Prior research has shown that mitigation strategies targeting a single demographic attribute – such as gender – can unintentionally intensify biases associated with other attributes, such as race [Ramachandranpillai et al.(2024)]. These unintended cross-demographic interactions may similarly apply to biases arising from perceptions of attractiveness. Moreover, existing approaches designed to counteract gender or race-based biases may inadvertently reinforce attractiveness related biases if these dimensions are not explicitly considered during model design and evaluation. Such inter-dependencies underscore the complexity of bias mitigation in MLLMs and highlight the need for further research into holistic, intersectional mitigation frameworks that address both traditional demographic variables and less-studied factors like attractiveness.
@@
Our work is not exempt from limitations. First, we do not evaluate API-based models such as GPT-4, as our experimental paradigm involves providing facial images to the models as inputs. Due to the lack of transparency related to data handling by API providers, and restrictions imposed by the dataset licenses (which specifically prohibits the use of images in such contexts) we excluded these models from our evaluation. Nevertheless, our analysis spans seven different open-source models, within which we observe statistically significant and consistent patterns, underscoring the robustness and prevalence of attractiveness-related biases. Second, while we carefully designed and curated our inputs to minimize confounding variables, our empirical evaluation may not fully capture the diversity of real-world contexts where attractiveness plays a role across cultures and social settings. Third, we have used a discrete number of categories for gender, age and race, as given by the ground truth labels of the dataset, which inevitably fails to capture and represent the diversity in society. Fourth, while it is possible to explore various prompt formulations, we limited our experiments to varying random seeds to ensure robustness. Our focus was on determining whether MLLMs respond differently to variations in facial attractiveness, an effect that our results consistently confirm. A systematic investigation of how this bias may be influenced by different prompting strategies remains an avenue for future research. Lastly, while we examine intersections with gender, age and race, we have not considered all relevant social factors, including socioeconomic status or disability, which could further impact the identified biases.
@@ 6 Conclusion
In this paper, we have studied the role that attractiveness plays in the decision-making processes of MLLMs by evaluating seven different models in 91 scenarios with over 900 face images and more than 7,000,000 prompts. Our findings provide strong evidence of the existence of an attractiveness bias in decisions made by MLLMs which can manifest in ways that mimic the attractiveness halo effect, a cognitive bias observed in humans. We also find evidence of the complex interplay between attractiveness and demographic factors –namely gender, age, and race– in driving the decisions of the MLLM. This interaction between attractiveness and other factors increases the complexity of interventions to mitigate biases in MLLMs. We hope that not only our results, but also our dataset and controlled experimental design, will serve as tools for the research community to further measure and understand these biases in MLLMs.
Acknowledgments
@@ Appendix C Scenarios
Model Makeup artist GPT-4
Acting professional Screen writer GPT-4
Personal trainer Engineer GPT-4
Real estate agent Computer Scientist GPT-4
Salesperson Product inspector GPT-4
News Anchor Copywriter GPT-4
Influencer Factory worker GPT-4
Public relations specialist Translator GPT-4
@@ Appendix E Attractiveness Halo Effect in Stereotyped Traits
Tables 10 - 16 below detail the strength of the attractiveness halo effect observed in all tested models. Each table reports the test statistic from the Kruskal–Wallis test (${H}_{i}^{\textit{attr}}$) comparing the values of $\phi_{i}$ between the beautified and original image groups for each scenario $s_{i}$. Standard star notation is employed to indicate the level of statistical significance: *** denotes $p<0.001$ and ** denotes $p<0.01$. Additionally, the tables provide the mean values of $p_{i}$ for both beautified and original images. Scenarios exhibiting statistically significant differences are highlighted in bold, with emphasis on the group displaying the higher mean value, indicating a stronger model tendency to associate that group with the first choice. This pattern reflects the model’s underlying preference or bias. Notably, across nearly all scenarios, the model demonstrates a consistent inclination to associate beautified images with the first choice, which corresponds to a positive sentiment (“Stereotyped Choice”). This trend underscores the presence of a robust attractiveness halo effect influencing the model’s decision-making.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 10: Attractiveness halo effect in sentiment oriented stereotyped traits for Gemma. Out of $3$ scenarios, 31 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 11: Attractiveness halo effect in sentiment oriented stereotyped traits for Phi3.5. Out of $3$ scenarios, 28 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 12: Attractiveness halo effect in sentiment oriented stereotyped traits for DeepSeek. Out of $3$ scenarios, 28 scenarios showed a significant attractiveness halo effect, while 3 showed an attractiveness bias but in the opposite direction.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 13: Attractiveness halo effect in sentiment oriented stereotyped traits for Molmo. Out of $3$ scenarios, 30 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 14: Attractiveness halo effect in sentiment oriented stereotyped traits for Qwen2. Out of $3$ scenarios, 28 scenarios showed a significant attractiveness halo effect, while 1 showed an attractiveness bias but in the opposite direction.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 15: Attractiveness halo effect in sentiment oriented stereotyped traits for Pixtral. Out of $3$ scenarios, 33 scenarios showed a significant attractiveness halo effect.
Choice 1 Choice 2 $\text{Mean }\phi_{i}(x^{b})$ $\text{Mean }\phi_{i}(x^{o})$ ${H}_{i}^{\textit{attr}}$
@@
Table 16: Attractiveness halo effect in sentiment oriented stereotyped traits for LLaVA 1.5. Out of $3$ scenarios, 32 scenarios showed a significant attractiveness halo effect.